Calculate the numerical aperture and acceptance angle for an optical diver having core of refractive index1.55 and cladding of refractive index1.50 respectively
1
Expert's answer
2020-04-15T10:34:23-0400

The refractive index of core, n1 = 1.55

and the refractive index of cladding , n2 = 1.50

Numerical aperture (NA) ="\sqrt{n_1^2-n_2^2}=\sqrt{1.55^2-1.50^2}=\sqrt{0.1525}=0.39"


Acceptance angle =sin-1(NA) = sin-1(0.39) = 23o